A ball is dropped from a hot air balloon moving vertically upward with a speed of 30m/s. If the balloon is 80m above the ground at the instant the ball isdropped. How long does it take for the ball to reach the ground?

Grade:12th pass

1 Answers

Arun
6 years ago
 
Firstly we have, t1=u/g or t1 = 30/10=3s 
H=u^2 /2g or 
H = 30*30/10=45
Secondly,
Total distance H=45+80=125
T=(2h/g)^1/2
T=5s 
 
Total time =t1+T=8s